The sum of the lengths of the two legs of right triangle ABC is 100 percent greater than the sum of the lengths of the legs of right triangle DEF, whose leg lengths are integers that sum to 7. If the two triangles are similar, what is the length of the hypotenuse of triangle ABC?
Answer by ewatrrr(24785)

100% greater than 7 is 14
The sum of the lengths of the two legs of right triangle ABC = 14
x^2 + (14-x)^2 = c^2
3^2 + 11^2 = 121, c = 11, the length of the hypotenuse of triangle ABC